# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v2

@@ 3.2.2. Implementation Strategies
[…]
Table 8. Container implementation strategies in multi-cloud environment
@@
Container Deployment and Management Deployment Strategies Automated Orchestration for Web Applications (Jha et al., 2019)
Kubernetes for Containerized APPs (Gattobigio et al., 2022)
Multi-cloud App Development Framework (Zhang et al., 2021)
Modular Approach to Container Implementation (Pham et al., 2015)
Deployment with Docker Images (Verdugo et al., 2017)
Docker Image Creation and Management (Mittal and Risco-Martín, 2017)
Real-time Kubernetes workloads in multi-cloud (Barletta et al., 2024)
AWS multi-cloud container app deployment (Kataru et al., 2023)
Multi-cloud container policy enforcement (Liu and Qin, 2024)
Probes and orchestration for cloud containers (Tundo et al., 2024)
Container Lifecycle Management Container Image Optimization (Al-shammari and Alwan, 2018)
Efficient Image Reorganization (Lorenzo et al., 2019)
Orchestrating with Docker and Kubernetes (Dimitrijevic et al., 2019)
Executable Images on Multi-cloud Platforms (Serhiienko and Spillner, 2018)
Privacy monitoring for containerized clouds (Barati et al., 2023)
Multi-cloud container resource usage monitoring (Horchulhack et al., 2024)
Container ecosystem for multi-cloud deployment (Wong et al., 2023)
Containers in AWS ECS for cloud services (Petrovski and Gusev, 2024)
Container management for multi-cloud MEC (Santos et al., 2023)
IoT container deployment in multi-cloud (Solayman and Qasha, 2023)
Data Management Firebase Cloud for Real-Time Database (Carpio et al., 2020)
High Availability with SpyStorage MCSS (Shen et al., 2017)
Data Storage with Object Store Service (Pop et al., 2016)
Container Orchestration Cloud Orchestration SLA-Based Container Strategies (Ramos et al., 2021)
Task Division and Scaling in Container Strategy (Bisht and Kaur, 2022)
Location-Aware Service Brokering (Chen et al., 2022)
Docker Containers for Multi-tenant Services (Slominski et al., 2015)
Resource Acquisition for Aneka Containers (Buyya and Barreto, 2015)
Cross-Cloud Container and VM Management (Janarthanan et al., 2018)
Container Orchestration Technologies Overview (Sami and Mourad, 2020)
Multi-cloud dynamic resource adaptation (Struhár et al., 2024)
AI/ML orchestration in multi-cloud containers (Sofia et al., 2024)
AI-driven container orchestration across clouds (Benmerar et al., 2023)
Cross-cloud container orchestration tools (Alonso et al., 2023a)
Kubernetes CI/CD for multi-cloud setups (Pandi et al., 2023)
Microservices Orchestration Microservices Architecture in IoT Cloud Apps (Ouyang et al., 2022)
ADAPT Deployment of Microservices (Alonso et al., 2019)
MSA Benchmark with Docker and Kubernetes (Rahman and Lama, 2019)
Kubernetes for power-efficient multi-cloud (Centofanti et al., 2024)
Edge computing with Docker and Kubernetes (Hossain et al., 2023)
Kubernetes for edge-cloud orchestration (Roy and Kalafatis, 2023)
Kubernetes orchestration for multi-cloud containers (Sofia et al., 2023)
Container Security Container Placement Secure Placement with Deep Learning (Deng et al., 2022)
Container Placement Strategies (Altran et al., 2022)
Dynamic microservice relocation across clouds (da Silva et al., 2023)
Security Policies SLA-Oriented Performance Isolation (Wang et al., 2020)
Access Control and Defense for Container Security (Gao et al., 2017)
HIP Integration for Docker Security (Ranjbar et al., 2017)
Secure Connectivity in Multi-tenant Environment (Alaluna et al., 2017)
Signed URLs for Access Control in Multi-cloud (Graupner et al., 2015a)
Performance and Scaling Strategies Performance Optimization LXC Virtual Cluster for Performance (Cui et al., 2016)
Scalable Multi-layer Architecture with Containers (Barna et al., 2017)
Reactive Auto-scaling and Resource Provisioning (Saif et al., 2023)
Istio for Control Plane Impact and Latency Reduction (Rodigari et al., 2021)
Auto-scaling containers across clouds (Dogani et al., 2023)
Multi-cloud resource monitoring and provisioning (Marques et al., 2024)
AI resource allocation in multi-cloud containers (Nagarajan et al., 2023)
@@
Scaling Strategies Edge Data Streaming and Real-Time Database (Carpio et al., 2020)
MicroCloud Architecture for Efficient Scaling (Baresi et al., 2016)
Distributed Workloads with Kubernetes (Lazuka et al., 2022)
Hybrid Abstractions for Container and Host (Bélair et al., 2019)
@@
Cloud Service and Networking Cloud Service Management Multi-Tenant Service with Docker and Cloudant (Slominski et al., 2015)
Abstraction Layer for Cloud Service Standardization (Pellegrini et al., 2017)
OpenStack and IoT in multi-cloud (Merlino et al., 2024)
Data indexing for multi-cloud exposure (Orzechowski et al., 2023)
Networking Edge Scheduling with Location-Awareness (Wong et al., 2019)
Openstack Networking with Kuryr (Bolivar et al., 2018)
Federated Framework for Cloud Storage (Harwalkar et al., 2019a)
Host Identity Protocol for Docker Networking (Ranjbar et al., 2017)
3.3. Pattern and Strategies (RQ2)
Table 9 presents various patterns and strategies employed in container-based applications in multi-cloud environment. These patterns and strategies are divided into various categories and subcategories, each of which highlights different aspects of multi-cloud architecture, management, security, resilience, and migration patterns and strategies. Notably, several studies (e.g., (Pahl, 2015; Graupner et al., 2015a; Shen et al., 2017; Pham and Pham, 2015)) appear across multiple subcategories because they report different patterns that address distinct aspects of cloud architecture and strategies. Every pattern is classified based on its specific focus, even when originating from the same study. Therefore, we categorized them into separate subcategories to reflect their unique contributions. Each category is briefly reported below.
@@ 3.3. Pattern and Strategies (RQ2)
2. Cloud Management and Resource Allocation Strategies: We identified 30 strategies within this category, grouped into four subcategories: Multi-Cloud Management Strategies, Container Management Strategies, Edge Computing and IoT Strategies, and Resource Management Strategies. The Multi-Cloud Management Strategies subcategory, with 8 strategies, focuses on managing and optimizing resources across multiple cloud environments. Leading strategies include the Multi-cloud Computing Strategy, Hybrid/Multi-cloud Approach, and MAPE-K control loop, all of which emphasize seamless integration and resource optimization across cloud platforms. The Container Management Strategies subcategory contains 7 strategies that address the management of containers in cloud environments. Key strategies include Docker Container Images and Lightweight Virtualization, both of which streamline container management and enhance resource efficiency in multi-cloud setups. The Edge Computing and IoT Strategies subcategory, with 8 strategies, highlights the integration of edge computing and IoT devices with cloud systems. Notable strategies include Mobile Edge Computing (MEC) and Fog Computing, which focus on distributing computation and storage closer to IoT devices for improved performance. Finally, the Resource Management Strategies subcategory includes 7 strategies that focus on efficient resource allocation in cloud environments. Leading approaches like Resource Allocation based on SLA levels, Container Orchestration, Kubernetes, Minikube, and AI-Driven Edge-Cloud Architecture emphasize intelligent resource distribution and orchestration across cloud and edge environments.
@@
Table 9. Patterns and strategies for container-based applications a multi-cloud environment
@@
Cloud Architectural Patterns and Models Architecture Patterns Microservice Architecture (Jha et al., 2019; Verdugo et al., 2017; Ouyang et al., 2022; Ranjbar et al., 2017; Quint and Kratzke, 2018; Pahl, 2015; Mittal and Risco-Martín, 2017; Sami and Mourad, 2020; Guerrero et al., 2018a; Centofanti et al., 2024; Hossain et al., 2023; Tiwari and Sharma, 2023)
Service-Oriented Architecture (SOA) (Buyya and Barreto, 2015; Casola et al., 2017; Gundu et al., 2020; Merlino et al., 2024)
Multitier Architecture, Layered Architecture (Jha et al., 2019; Ramos et al., 2021; Bucur et al., 2018; Barati et al., 2023; Horchulhack et al., 2024; Liu and Qin, 2024; Tundo et al., 2024)
Client-Server Architecture (Hua et al., 2020; Bucur et al., 2018)
Component-based Architecture (Wong et al., 2019)
Event-driven Architecture (Wurster et al., 2018)
Multi-agent Architecture (Mfula and Nurminen, 2018; Saif et al., 2023)
Master-worker Nodes Architecture (Sami and Mourad, 2020)
Netflix Zuul-based Seeker Component (Slominski et al., 2015)
Multi-tenant Cloud Service Architecture (Slominski et al., 2015)
Stateful Engine Architecture (Slominski et al., 2015)
Distributed Multi-Cloud Native Architecture (Alonso et al., 2023a)
Plugin-Based Deployment Automation Architecture (Harzenetter et al., 2023)
Cloud-Edge Integrated Robotics Architecture (Huang and Wang, 2023)
Hybrid Malware Detection Architecture (Wang et al., 2023)
Cloudfront-Enhanced Container Architecture (Kataru et al., 2023)
AI-Driven Edge-Cloud Architecture (Benmerar et al., 2023)
Metadata-Driven Architecture (Orzechowski et al., 2023)
Blockchain-Based Architecture (Ahmad and Aujla, 2023)
Cross-Layered Edge-Cloud Architecture (Sofia et al., 2023; Santos et al., 2023)
Communication and Networking Patterns Service Mesh (Gattobigio et al., 2022)
Service Chaining (SC) (Bhamare et al., 2015)
Multi-Cloud APIs (Graupner et al., 2015a)
Third-Party APIs (Graupner et al., 2015a)
Service Request Broker (Zhang et al., 2021)
Publish and Subscribe Communication Protocol (Trihinas et al., 2015)
Synchronous Network Communication Protocols (Almeida et al., 2014)
Asynchronous Network Communication Protocols (Almeida et al., 2014)
Network Function Virtualization (NFV) (Bhamare et al., 2015)
Service Discovery (Pop et al., 2016)
Blockchain-Based Identity and Access Management (IAM) Architecture (Fugkeaw, 2023)
Federated Cloud-Edge Architecture (Roy and Kalafatis, 2023)
Deployment Pattern black-Green deployment (Alonso et al., 2019; Sheridan et al., 2017)
Object store service (Pop et al., 2016)
Multi-cloud Deployment Model (Sabbioni et al., 2020)
Distributed Deployment Model (Sabbioni et al., 2020)
Multi-Cloud Deployment Architecture (Petrovski and Gusev, 2024; Pandi et al., 2023)
Decentralized Orchestration Architecture (Sofia et al., 2024)
Policy-Driven Orchestration Architecture (Lee and Nam, 2023)
DevOps-Oriented Architecture (Solayman and Qasha, 2023)
Service Models Platform-as-a-Service (PaaS) (Pahl, 2015; Shi et al., 2020)
Software-as-a-Service (Serhiienko and Spillner, 2018)
Infrastructure-as-a-Service (IaaS) (Pahl, 2015)
Function-as-a-Service (FaaS) (Pahl, 2015; Ramesh et al., 2023)
Multi-Cloud Orchestration Architecture (Pandi et al., 2023)
Serverless Container-Oriented Architecture (Shi et al., 2020)
@@
Hybrid cloud architecture (Wang et al., 2019)
Hybrid/Multi-cloud Approach (Gattobigio et al., 2022; Wahab et al., 2016)
Introducing multi-cloud Middleware (Casola et al., 2017; Acquaviva et al., 2017)
MAPE-K control loop (Fadda et al., 2019; Saif et al., 2023)
Connecting to multiple cloud service providers (Pop et al., 2016)
Multi-cloud load balancing (Wang et al., 2020; Pop et al., 2016)
Multi-Cloud BP provisioning (Kritikos et al., 2019)
Cross-level orchestration of cloud services (Kritikos et al., 2019)
Cross-level monitoring and adaptation of BPs (Kritikos et al., 2019)
Container Management Strategies Linux Container (LXC) project (Pahl, 2015)
Container engine (Pahl, 2015)
Docker Container Images (Pahl, 2015)
Lightweight virtualization (Pahl, 2015)
Portable application packaging (Pahl, 2015)
One-container-per-app approach (Pahl, 2015)
Data volumes and data volume containers (Pahl, 2015)
Edge Computing and IoT Strategies Mobile Edge Computing (MEC) (Wong et al., 2019; Lazuka et al., 2022; Dreibholz et al., 2019; Bucur et al., 2018)
Fog computing (Al-shammari and Alwan, 2018)
Edge Services (Carpio et al., 2020)
Connecting IoT Edge Devices to the Cluster (Gattobigio et al., 2022)
Cloud of Things (CoT) (Gattobigio et al., 2022)
Multi-Access Edge Computing (MEC) Architecture (Santos et al., 2023)
Cross-Layered Edge-Cloud Architecture (Sofia et al., 2023)
Cloud-Edge Integrated Robotics Architecture (Huang and Wang, 2023)
AI-Driven Edge-Cloud Architecture (Benmerar et al., 2023)
Resource Management Strategies Resource allocation based on different SLA levels (Wang et al., 2020)
Container Orchestration, Kubernetes, Minikube (Pahl, 2015; Gupta et al., 2021; Sami and Mourad, 2020)
Rate-Based Stream Processing Architecture (Nguyen and Chien, 2023)
Multi-Agent Resource Optimization Architecture (Nagarajan et al., 2023)
@@
Cloud Security and Resilience Strategies Security and Resiliency Patterns Security-by-Design approach (Buyya and Barreto, 2015)
Encryption of files (Graupner et al., 2015a)
Standardized APIs for file transfer (Graupner et al., 2015a)
Centralized/External ACM Service (Graupner et al., 2015a)
Distributed ACM Service (Graupner et al., 2015a)
Dynamic Switching between Authentication Methods (Graupner et al., 2015a)
ACM based on Signed URLs (Graupner et al., 2015a)
Attribute-Based Encryption and Signature (Shen et al., 2017)
Secure data sharing (Shen et al., 2017)
Data access control (Shen et al., 2017)
Local encryption and signing (Shen et al., 2017)
Authorized Tokens (Shen et al., 2017)
Byzantine quorum protocol (Shen et al., 2017)
Blockchain-Driven SLA Management Architecture (Neeraj et al., 2023)
Hybrid Malware Detection Architecture (Wang et al., 2023)
Fault-tolerance Strategies Fault-tolerance with Redundant Engines (Slominski et al., 2015)
Multi-cloud systems fault-tolerant workflow (Tang, 2021)
RAFT Consensus Algorithm (Naik, 2016a)
Cloud Migration Strategies Service-oriented Migration (Pham and Pham, 2015)
Application-centric Migration (Pham and Pham, 2015)
Image-based Migration (Pham and Pham, 2015)
Migration to a Virtualized Container (Pham and Pham, 2015)
[…]
3.4. Quality Attributes and Tactics (RQ3)
Table 10 provides a comprehensive overview of the quality attributes and related tactics for containerized applications in multi-cloud environment. It is based on data extracted from 121 selected studies, which are organized following the ISO 25010 standard. This standard defines high-level QAs along with their sub-characteristics. Table 10 lists nine QAs, namely Performance (Efficiency), Security, Compatibility, Scalability, Reliability, Portability, Flexibility, Maintainability, and Usability. Each QA has its related terms, associated study numbers, and corresponding tactics to achieve them. In Table 10, due to limited space, we provide only five tactics for each QA. The complete set of 70 tactics is listed in the provided dataset (Waseem et al., 2024). However, it is important to note that these tactics are scattered across different studies. Below we briefly discuss how each QA can be achieved by applying the mentioned tactics.

### v3

@@ 3.2.2 Implementation Strategies
Table 8: Container implementation strategies in multi-cloud environment
@@
Container Deployment and Management Deployment Strategies Automated Orchestration for Web Applications [96]
Kubernetes for Containerized APPs [34]
Multi-cloud App Development Framework [101]
Modular Approach to Container Implementation [86]
Deployment with Docker Images [120]
Docker Image Creation and Management [106]
Real-time Kubernetes workloads in multi-cloud [82]
AWS multi-cloud container app deployment [134]
Multi-cloud container policy enforcement [130]
Probes and orchestration for cloud containers [91]
Container Lifecycle Management Container Image Optimization [112]
Efficient Image Reorganization [113]
Orchestrating with Docker and Kubernetes [61]
Executable Images on Multi-cloud Platforms [45]
Privacy monitoring for containerized clouds [135]
Multi-cloud container resource usage monitoring [87]
Container ecosystem for multi-cloud deployment [136]
Containers in AWS ECS for cloud services [100]
Container management for multi-cloud MEC [125]
IoT container deployment in multi-cloud [38]
Data Management Firebase Cloud for Real-Time Database [30]
High Availability with SpyStorage MCSS [109]
Data Storage with Object Store Service [32]
Container Orchestration Cloud Orchestration SLA-Based Container Strategies [132]
Task Division and Scaling in Container Strategy [39]
Location-Aware Service Brokering [102]
Docker Containers for Multi-tenant Services [44]
Resource Acquisition for Aneka Containers [98]
Cross-Cloud Container and VM Management [123]
Container Orchestration Technologies Overview [29]
Multi-cloud dynamic resource adaptation [137]
AI/ML orchestration in multi-cloud containers [138]
AI-driven container orchestration across clouds [104]
Cross-cloud container orchestration tools [33]
Kubernetes CI/CD for multi-cloud setups [117]
Microservices Orchestration Microservices Architecture in IoT Cloud Apps [121]
ADAPT Deployment of Microservices [105]
MSA Benchmark with Docker and Kubernetes [53]
Kubernetes for power-efficient multi-cloud [139]
Edge computing with Docker and Kubernetes [37]
Kubernetes for edge-cloud orchestration [46]
Kubernetes orchestration for multi-cloud containers [118]
Container Security Container Placement Secure Placement with Deep Learning [69]
Container Placement Strategies [114]
Dynamic microservice relocation across clouds [115]
Security Policies SLA-Oriented Performance Isolation [66]
Access Control and Defense for Container Security [70]
HIP Integration for Docker Security [68]
Secure Connectivity in Multi-tenant Environment [124]
Signed URLs for Access Control in Multi-cloud [57]
Performance and Scaling Strategies Performance Optimization LXC Virtual Cluster for Performance [126]
Scalable Multi-layer Architecture with Containers [52]
Reactive Auto-scaling and Resource Provisioning [77]
Istio for Control Plane Impact and Latency Reduction [35]
Auto-scaling containers across clouds [116]
Multi-cloud resource monitoring and provisioning [88]
AI resource allocation in multi-cloud containers [43]
@@
Scaling Strategies Edge Data Streaming and Real-Time Database [30]
MicroCloud Architecture for Efficient Scaling [31]
Distributed Workloads with Kubernetes [74]
Hybrid Abstractions for Container and Host [71]
@@
Cloud Service and Networking Cloud Service Management Multi-Tenant Service with Docker and Cloudant [44]
Abstraction Layer for Cloud Service Standardization [73]
OpenStack and IoT in multi-cloud [36]
Data indexing for multi-cloud exposure [110]
Networking Edge Scheduling with Location-Awareness [111]
Openstack Networking with Kuryr [40]
Federated Framework for Cloud Storage [108]
Host Identity Protocol for Docker Networking [68]
Takeaway 5: The reviewed studies identify a wide range of implementation strategies that reflect the specific challenges of containerizing applications in multi-cloud environments. These include orchestrating containers across heterogeneous platforms, ensuring secure placement and access control under varying provider policies, and optimizing performance and scaling through AI-driven resource provisioning and hybrid abstractions. Unique to multi-cloud settings are strategies for managing distributed services, networking across clouds, and integrating edge and IoT services.
3.3 Pattern and Strategies (RQ2)
Table 9 presents a thematic classification of patterns and strategies employed in container-based applications within multi-cloud environments. These patterns were identified through qualitative coding and grouped into multiple categories and subcategories based on their architectural, management, security, resilience, and migration focus. Each row corresponds to a pattern extracted from the literature, and studies that reported multiple distinct patterns appear in more than one subcategory (e.g., [60, 57, 109, 80]). Table 9 is structured to allow readers to quickly identify and compare solution strategies across different technical domains. To improve interpretability, we provide a brief overview of each category below.
@@ 3.3 Pattern and Strategies (RQ2)
2. Cloud Management and Resource Allocation Strategies: We identified 30 strategies within this category, grouped into four subcategories: Multi-Cloud Management Strategies, Container Management Strategies, Edge Computing and IoT Strategies, and Resource Management Strategies. The Multi-Cloud Management Strategies subcategory, with 8 strategies, focuses on managing and optimizing resources across multiple cloud environments. Leading strategies include the Multi-cloud Computing Strategy, Hybrid/Multi-cloud Approach, and MAPE-K control loop, all of which emphasize seamless integration and resource optimization across cloud platforms. The Container Management Strategies subcategory includes 7 approaches that focus on effectively managing containers in cloud environments. Key examples are Docker Container Images and Lightweight Virtualization, both of which help simplify container operations and improve resource efficiency in multi-cloud settings. The Edge Computing and IoT Strategies subcategory consists of 8 strategies that explore how edge computing and IoT devices can be integrated with cloud systems. Noteworthy strategies such as Mobile Edge Computing (MEC) and Fog Computing aim to bring computation and storage closer to IoT devices, enhancing system responsiveness and performance. Lastly, the Resource Management Strategies subcategory comprises 7 strategies dedicated to optimizing resource usage in cloud environments. Prominent approaches like Resource Allocation based on SLA levels, Container Orchestration, Kubernetes, Minikube, and AI-Driven Edge-Cloud Architecture focus on smart resource distribution and coordination across cloud and edge platforms.
@@
Takeaway 6: The studies reveal that when containerized applications are deployed in multi-cloud environments, architecture and management patterns must be adapted to support distributed, heterogeneous infrastructure. Patterns like Microservices and SOA are extended with multi-cloud orchestration and federation capabilities, while service mesh and chaining are tailored for secure, cross-cloud communication. Deployment and fault-tolerant patterns—such as black-Green deployment, decentralized orchestration, and blockchain-driven resilience.
Table 9: Patterns and strategies for container-based applications in multi-cloud environment
@@
Cloud Architectural Patterns and Models Architecture Patterns Microservice Architecture [96, 120, 121, 68, 50, 60, 106, 29, 99, 139, 37, 131]
Service-Oriented Architecture (SOA) [98, 127, 64, 36]
Multitier Architecture, Layered Architecture [96, 132, 133, 135, 87, 130, 91]
Client-Server Architecture [93, 133]
Component-based Architecture [111]
Event-driven Architecture [140]
Multi-agent Architecture [42, 77]
Master-worker Nodes Architecture [29]
Netflix Zuul-based Seeker Component [44]
Multi-tenant Cloud Service Architecture [44]
Stateful Engine Architecture [44]
Distributed Multi-Cloud Native Architecture [33]
Plugin-Based Deployment Automation Architecture [141]
Cloud-Edge Integrated Robotics Architecture [49]
Hybrid Malware Detection Architecture [142]
Cloudfront-Enhanced Container Architecture [134]
AI-Driven Edge-Cloud Architecture [104]
Metadata-Driven Architecture [110]
Blockchain-Based Architecture [143]
Cross-Layered Edge-Cloud Architecture [118, 125]
Communication and Networking Patterns Service Mesh [34]
Service Chaining (SC) [47]
Multi-Cloud APIs [57]
Third-Party APIs [57]
Service Request Broker [101]
Publish and Subscribe Communication Protocol [84]
Synchronous Network Communication Protocols [95]
Asynchronous Network Communication Protocols [95]
Network Function Virtualization (NFV) [47]
Service Discovery [32]
Blockchain-Based Identity and Access Management (IAM) Architecture [144]
Federated Cloud-Edge Architecture [46]
Deployment Pattern black-Green deployment [105, 55]
Object store service [32]
Multi-cloud Deployment Model [65]
Distributed Deployment Model [65]
Multi-Cloud Deployment Architecture [100, 117]
Decentralized Orchestration Architecture [138]
Policy-Driven Orchestration Architecture [145]
DevOps-Oriented Architecture [38]
Service Models Platform-as-a-Service (PaaS) [60, 97]
Software-as-a-Service [45]
Infrastructure-as-a-Service (IaaS) [60]
Function-as-a-Service (FaaS) [60, 92]
Multi-Cloud Orchestration Architecture [117]
Serverless Container-Oriented Architecture [97]
@@
Hybrid cloud architecture [75]
Hybrid/Multi-cloud Approach [34, 72]
Introducing multi-cloud Middleware [127, 146]
MAPE-K control loop [85, 77]
Connecting to multiple cloud service providers [32]
Multi-cloud load balancing [66, 32]
Multi-Cloud BP provisioning [76]
Cross-level orchestration of cloud services [76]
Cross-level monitoring and adaptation of BPs [76]
Container Management Strategies Linux Container (LXC) project [60]
Container engine [60]
Docker Container Images [60]
Lightweight virtualization [60]
Portable application packaging [60]
One-container-per-app approach [60]
Data volumes and data volume containers [60]
Edge Computing and IoT Strategies Mobile Edge Computing (MEC) [111, 74, 122, 133]
Fog computing [112]
Edge Services [30]
Connecting IoT Edge Devices to the Cluster [34]
Cloud of Things (CoT) [34]
Multi-Access Edge Computing (MEC) Architecture [125]
Cross-Layered Edge-Cloud Architecture [118]
Cloud-Edge Integrated Robotics Architecture [49]
AI-Driven Edge-Cloud Architecture [104]
Resource Management Strategies Resource allocation based on different SLA levels [66]
Container Orchestration, Kubernetes, Minikube [60, 103, 29]
Rate-Based Stream Processing Architecture [119]
Multi-Agent Resource Optimization Architecture [43]
@@
Cloud Security and Resilience Strategies Security and Resiliency Patterns Security-by-Design approach [98]
Encryption of files [57]
Standardized APIs for file transfer [57]
Centralized/External ACM Service [57]
Distributed ACM Service [57]
Dynamic Switching between Authentication Methods [57]
ACM based on Signed URLs [57]
Attribute-Based Encryption and Signature [109]
Secure data sharing [109]
Data access control [109]
Local encryption and signing [109]
Authorized Tokens [109]
Byzantine quorum protocol [109]
Blockchain-Driven SLA Management Architecture [147]
Hybrid Malware Detection Architecture [142]
Fault-tolerance Strategies Fault-tolerance with Redundant Engines [44]
Multi-cloud systems fault-tolerant workflow [41]
RAFT Consensus Algorithm [148]
Cloud Migration Strategies Service-oriented Migration [80]
Application-centric Migration [80]
Image-based Migration [80]
Migration to a Virtualized Container [80]
3.4 Quality Attributes and Tactics (RQ3)
Table 10 presents a thematic classification of QAs and their associated implementation tactics for containerized applications in multi-cloud environments. These QAs and tactics were derived through thematic analysis of data extracted from 121 primary studies and are organized according to the ISO 25010 standard. This standard defines high-level software quality characteristics and their sub-attributes, which served as a guiding framework for our classification. The table includes nine core QAs, namely Performance (Efficiency), Security, Compatibility, Scalability, Reliability, Portability, Flexibility, Maintainability, and Usability. For each QA, the table lists representative related terms, example study references, and commonly applied tactics. Due to space constraints, only five example tactics per QA are shown in Table 10, while the complete list of 70 tactics is available in the replication package [11]. These tactics are distributed across multiple studies and highlight recurring practices for achieving specific quality goals. To help readers navigate this classification, we briefly explain below how each QA is addressed in the literature and what patterns emerge in terms of implementation strategies.